# Higgs Inflation Model with Small Non-Minimal Coupling Constant

### v4

@@ Abstract
In the proposed model of the "Higgs field$+$gravity" system, the self-consistency of the system of equations obtained from the original action has the form of an algebraic equation (constraint) defining the scalar $\zeta$ as a function of the Higgs field $\varphi$ and its first derivatives. The scalar $\zeta$ is present in all equations of motion and has a significant effect on the dynamics of the fields. After the transition in the equations of motion to the Einstein frame with the spatially flat Friedmann metric, it is convenient to describe the resulting system of equations using the action $S_{eff}$ and the Lagrangian $L_{eff}$, which we call the effective action and the effective Lagrangian and from which these equations can be obtained. Due to the constraint, the original model parameters are converted in $L_{eff}$ into $\varphi$-dependent classical effective parameters. In particular, the effective potential $U_{eff}(\varphi)$ in $L_{eff}$ has the form $U_{eff}=\frac{\lambda}{4\xi^{2}}M_{P}^{4}\cdot F(\varphi)\cdot\tanh^{4}\bigl{(}\frac{\sqrt{\xi}\varphi}{M_{P}}\bigr{)}$, where $F(\varphi)$ is a smooth function equal to $F(\varphi)\approx\frac{1}{2}$ for $\varphi\gg\sqrt{3/8}M_{P}$. The constant $\xi$ of non-minimal coupling to scalar curvature can be chosen as small as desired. If $\xi=1/6$, then to ensure agreement with CMB observational data, the Higgs field self-coupling parameter $\lambda$ in the original action must be of the order of $\sim 10^{-11}$. During cosmological evolution after the end of inflation, the decrease of $\varphi$ leads to a change in the sign of the effective Higgs mass term in $L_{eff}$. This TMSM effect provides an answer to the mystery of the Higgs potential structure and leads to spontaneous symmetry breaking. As $\varphi$ approaches VEV, the scalar function $\zeta(\varphi)$ changes in such a way that the classical effective self-coupling parameter $\lambda_{eff}(\zeta(\varphi))$ increases by 10 orders of magnitude compared to $\lambda$, which is necessary for the implementation of the GWS theory. Applying the model to the very beginning of the classical evolution of the Universe shows that under certain initial conditions, cosmological dynamics can begin with a "pathological" and even phantom regime preceding inflation. However, if evolution begins with normal dynamics, then it proceeds only as inflation, and the problem of initial conditions for the onset of inflation does not arise.
@@ 1 Introduction
To compare all existing Higgs inflation models with the model proposed in this paper, we begin with a brief description of the main steps of their construction at the tree level. The electroweak SM at the tree level is defined in Minkowski space. To use the SM in cosmology there is a need to reformulate it in a general covariant form to allow gravity to be included in the model. The next steps are to add a nonminimal coupling of the Higgs field to the scalar curvature, followed by the transition to the Einstein frame in the original action and a redefinition of the Higgs field leading to a Lagrangian with a canonically normalized field. This combination of particle physics and gravity, which we will call SM+Gravity, is the basis for how the Higgs sector of the electroweak SM is commonly used to study cosmology at energies well above the electroweak energy scale. However, the resulting models are suitable for describing inflation in accordance with the CMB data [20], [21] only if the constant of non-minimal coupling to scalar curvature is unnaturally huge: depending on the type of model, metric or Palatini, the coupling constant varies in the range $10^{4}\lesssim\xi\lesssim 10^{8}$. Nevertheless, apart from some effects due to nonminimal coupling (see, for example, review [10]), in the SM+Gravity approach to Higgs inflation models at the tree level, the question of finding a reasonable connection between low-energy particle physics and physics at the inflation energy scale remains open. Moreover, the need to take into account quantum corrections runs into problems associated with the non-renormalizability of the resulting models, which further increases doubts about the validity of the SM+Gravity approach.
In the present paper, the Higgs inflation model is formulated in the framework of the Two Measure Theory (TMT), which radically changes the approach to combining gravity with particle physics. Due to these changes, a formulation of the SM that takes into account the structure and procedure of TMT (and which we will refer to as Two-Measure Standard Model (TMSM)) can successfully overcome the above-mentioned problems of Higgs inflation.
@@
  • The main result of the present paper is the demonstration that the Higgs+gravity sector of TMSM with an arbitrarily small constant $\xi$ of non-minimal coupling to the scalar curvature allows one to construct a model of Higgs inflation that agrees well with observational CMB data. At the same time, despite the seemingly obvious expectations, there is no contradiction between this result and the fact that the TMSM also gives the correct values of the SM physical quantities at accelerator energies. Moreover, the TMSM provides a natural explanation for the mystery of the appearance of a mass term with the "wrong" sign in the Higgs potential as an effect arising during cosmological evolution due to a special property of the TMSM associated with the presence of two volume measures in the original action.
@@
are well known as the constraints neeeded for the onset of inflation. The initial understanding of this issue was developed in the first models of chaotic inflation [23], [24]. However, later, the need to implement flat potential models [25], [26], [27], [28], [29], [2], in which inflation predictions agree well with the SMB observational data, turned the question of initial conditions into a problem. The seriousness and urgency of this problem became clear from the discussion in papers [30]-[34] . A completely new approach to the problem of initial conditions for inflation based on the TMT was proposed in ref. [35], where it was shown that the solution to the problem is closely related to the need to answer the question of what happened before inflation. Within the framework of the TMT model of ref. [35], where the inflaton field was considered independent of the known matter fields, the problem of initial conditions for inflation turned out to be related to the GBV theorem [36]-[39]. In the present paper, in which the inflaton field is identified with the Higgs field, we find that not only are the initial conditions (1.1) for inflation guaranteed to be satisfied, but an earlier, rather exotic stage of evolution preceding inflation is also possible.
@@
The organization of the paper is as follows. In section 2, after defining the primordial action [^2] In ”conventional” alternative gravity theories, the original action differs from the Einstein-Hilbert action only in the form of the Lagrangian, for example, by the presence of a non-minimal coupling with curvature. The transformation to the Einstein frame, which in conventional theories is performed in the original action, simply changes the set of variables used to describe the theory, but does not change the theory itself. In TMT, the main difference is that the volume measure contains $\Upsilon$, although a modification of the Lagrangian is also possible. Therefore, the original TMT action differs from the Einstein-Hilbert action in the form of the Lagrangian density, and to bring the TMT action to the Einstein frame, the conformal transformation must exclude from the gravitational term of the action not only the non-minimal coupling, but also $\Upsilon$. In a theory obtained in this way, the constraint determining $\zeta$ cannot arise, that is, by acting in this way we would be dealing with a different theory. It is therefore incorrect to use the term ”Jordan frame” in the usual sense for the set of variables in the original action of TMT. To emphasize this distinction, we will use the terms ”primordial variables”, ”primordial model parameters”, ”primordial Lagrangian” and ”primordial action” instead of ”variables in the original frame”, ”original model parameters”, ”original Lagrangian” and ”original action”. and justifying the need to add a new type of vacuum-like term related to the volume measure density $\Upsilon$, we perform all the steps required by the TMT procedure. Having found the equations and represented them in the Einstein frame in $\zeta$-dependent form, we, using the constraint, arrive at the TMT-effective equations, potential, Lagrangian and action [^3] Note that this use of the term ”effective” refers to the classical description and, for example, is in no way connected with the usual use of the term ”effective potential” calculated taking into account quantum corrections. for the Higgs sector of the TMSM (where all gauge field and fermion contributions are omitted). Their analysis shows that the obtained (classical) TMT-effective action is applicable in the entire admissible energy spectrum: from the energies of the electroweak SM to the scales of inflation. In section 3, after choosing the model parameters that provide the required flatness of the TMT effective potential for values of the canonically normalized scalar field $\varphi>6M_{P}$, we study a simplified model showing its agreement with the observational CMB data. The need to consider a simplified model (which will be justified in section 6) is due to the fact that the found TMT-effective action has a structure reminiscent of the K-essence models commonly used to describe the late time cosmic acceleration [40]-[43]. Important details of the shape of the TMT-effective potential and the behavior of the function $\zeta(\varphi)$ during the entire cosmological evolution are analyzed in section 4. This allows us to understand the role of cosmological evolution in the appearance of the negative mass term in the Higgs potential. Section 5, which is partly related to the paper [35], focuses on the effects dictated by the need to ensure that cosmological solutions (and their initial conditions) must keep the $\zeta$ function positive. In section 6 we study the influence of the K-essence type structure on the earliest stages of classical evolution of the Universe. We show that the K-essence type structure may be responsible for the existence of the earliest stages of evolution, which have "pathological" dynamics. However, if evolution starts with "normal" dynamics, then it proceeds only as inflation, and the problem of initial conditions for the onset of inflation does not arise. Finally, it is shown that the K-essence type structure does not change the results of the simplified version of the model obtained in section 3, as far as the slow-roll inflation stage is concerned. However, the K-essence type structure may have some observable effect in the late stage of inflation. In section 7 we focus on discussing two fundamentally important features of TMSM that provide the main results of the paper.
@@ 2.1 Action
In this paper we will study the TMSM with the following primordial action
$S_{primordial}=\int d^{4}x(\sqrt{-g}+\Upsilon)\left(-\frac{M_{P}^{2}}{2}\right)\Biggl{(}1+\xi\frac{2|H|^{2}}{M_{P}^{2}}\Biggr{)}R(\Gamma,g)+S_{H}+S_{vac}$ (2.1)
Here $M_{P}$ is the reduced Planck mass. We will proceed in Palatini’s formulation, where $\Gamma$ denotes the affine connection, and $R(\Gamma,g)=g^{\mu\nu}R_{\mu\nu}(\Gamma)$, $R_{\mu\nu}(\Gamma)=R^{\lambda}_{\mu\nu\lambda}(\Gamma)$ and $R^{\lambda}_{\mu\nu\sigma}(\Gamma)\equiv\Gamma^{\lambda}_{\mu\nu,\sigma}+\Gamma^{\lambda}_{\gamma\sigma}\Gamma^{\gamma}_{\mu\nu}-(\nu\leftrightarrow\sigma)$. $H$ is the Higgs isodoublet. Following what was said in the Introduction, all contributions from gauge fields and fermions are omitted. In our notations, in a theory with only the volume element $\sqrt{-g}d^{4}x$, the model of massless scalar field non-minimally coupled to scalar curvature would be conformally invariant if the parameter $\xi$ were equal to $\xi=-\frac{1}{6}$. The primordial action for the Higgs field $H$ is chosen as follows
$S_{H}=\int d^{4}x\left[(b_{k}\sqrt{-g}-\Upsilon)g^{\alpha\beta}\left(\partial_{\alpha}H\right)^{{\dagger}}\partial_{\beta}H-(b_{p}\sqrt{-g}+\Upsilon)\lambda|H|^{4}-(b_{p}\sqrt{-g}-\Upsilon)m^{2}|H|^{2}\right],$ (2.2)
The model parameters $b_{k}>0$, $b_{p}>0$, $\lambda>0$ and $m^{2}>0$ are chosen to be positive. Given that the FLRW Universe is a model resulting from cosmological averaging, we can consider the corresponding average of $H$. Let $\phi/\sqrt{2}$ denote the only non-zero component of $H$ obtained by cosmological averaging, and we assume that $\phi$ plays the role of the inflaton.
Contribution of the vacuum-like terms to the primordial action is defined by
@@
If the first term in eq.(2.3) was present in Einstein’s GR, $V_{1}$ would be a cosmological constant. The term with $V_{2}$ was first introduced in Ref. [46]. The first reason for adding the term with $V_{2}$ is that the term $\propto\Upsilon=\zeta(x)\sqrt{-g}$, which can be expected as the contribution of quantum-gravitational effects to a vacuum-like action, does not contribute to the equations of motion. Therefore, the next in powers of $\zeta$ vacuum-like term can be of the form $\propto\zeta^{2}\sqrt{-g}=\Upsilon^{2}/\sqrt{-g}$. There are also more pragmatic reasons. It turns out that thanks to this term, the corresponding TMT-effective potential and the TMT-effective action acquire fundamentally new important properties without which the goals of this paper could not be achieved.
@@ 2.2 Equations of motion in the Einstein frame
$-\frac{M_{P}^{2}}{2}\Omega R(\Gamma,g)-\frac{1}{2}g^{\alpha\beta}\phi_{,\alpha}\phi_{,\beta}-\frac{\lambda}{4}\phi^{4}+\frac{m^{2}}{2}\phi^{2}-2\zeta V_{2}={\mathcal{M}}$ (2.8)
@@
$(1-\zeta)\left(-\frac{M_{P}^{2}}{2}\right)\Omega R_{\mu\nu}(\Gamma)+(b_{k}-\zeta)\frac{1}{2}\phi_{,\mu}\phi_{,\nu}$
@@
Using the trace of eq.(2.9) one can eliminate the term $-\frac{1}{2}M_{P}^{2}\Omega R(\Gamma,g)$ from eq.(2.8). As a result, we find that for these equations to be consistent, it is necessary that the scalar function $\zeta(x)$ satisfy the following relation
@@
with the same Newton constant as in the original frame. Here $R_{\mu\nu}(\tilde{g})$ and $R(\tilde{g})$ are the Ricci tensor and the scalar curvature of the metric $\tilde{g}_{\mu\nu}$, respectively. Therefore, gravity becomes canonical, and the set of dynamical variables using the metric $\tilde{g}_{\mu\nu}$ can be called the Einstein frame. $T_{\mu\nu}^{(eff)}$ on the right side of the Einstein equations (2.14) is the TMT-effective energy-momentum tensor
@@
$\zeta=\zeta(\phi,X_{\phi};\mathcal{M})=\frac{\mathcal{M}-2V_{1}-(2b_{p}-1)\frac{\lambda}{4}\phi^{4}-(2b_{p}+1)\frac{m^{2}}{2}\phi^{2}+(1+b_{k})\Omega X_{\phi}}{\mathcal{M}-2V_{2}+\frac{\lambda}{4}\phi^{4}-\frac{m^{2}}{2}\phi^{2}+(1+b_{k})\Omega X_{\phi}}.$ (2.17)
@@
In the Higgs inflation models in the conventional theory (both in the metric and in the Palatini formulation), the transition to the Einstein frame performed directly in the original action completely eliminates the non-minimal coupling of the scalar field to the scalar curvature. In the Two-Measure Standard Model under study, as in all TMT models, the transition to the Einstein frame (2.13) is carried out in the equations of motion and we notice that in the $\phi$-field equation (2.18) presented in the Einstein frame using (2.13), the term of the non-minimal coupling $\propto\xi R(\tilde{g})\phi$ is preserved. After passing to the Einstein frame in eq.(2.18) using (2.13), we can insert the expression for the scalar curvature $R(\tilde{g})$ obtained from the Einstein equations (2.14). Then, the $\phi$-field equation in the Einstein frame takes the following form
$\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}\Bigg{(}\frac{b_{k}-\zeta}{(1+\zeta)\Omega}\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}\phi\Bigg{)}+\frac{b_{p}+\zeta}{(1+\zeta)^{2}\Omega^{2}}\lambda\phi^{3}+\frac{b_{p}-\zeta}{(1+\zeta)^{2}\Omega^{2}}m^{2}\phi$
$+\frac{2\xi\phi}{(1+\zeta)M_{P}^{2}\Omega^{3}}\left[-\mathcal{M}-2\zeta V_{2}-\frac{\lambda}{4}\phi^{4}+\frac{m^{2}}{2}\phi^{2}-(1+\zeta)\Omega X_{\phi}\right]=0.$ (2.19)
The procedure performed means that we are dealing with an explicitly formulated self-consistent system of the Higgs field and gravity.

### v5

@@ Abstract
The Higgs sector of the Two-Measure Theory (TMT) extension of the electroweak SM (TMSM) is studied in the context of cosmology, where the only non-zero component $\varphi$ of the cosmologically averaged Higgs field plays the role of the inflaton. The self-consistency of the system of equations obtained from the original action has the form of an algebraic constraint defining the scalar $\zeta$, which is the ratio of two volume measures, as a function of the field $\varphi$ and its first derivatives. The scalar $\zeta$ is present in all equations of motion and has a significant effect on the dynamics of the fields. After the transition in the equations of motion to the Einstein frame with the spatially flat Friedmann metric, it is convenient to describe the resulting system of equations using the action $S_{eff}$ and the Lagrangian $L_{eff}$, which we call the TMT-effective action and the TMT-effective Lagrangian and from which these equations can be obtained. Due to the constraint, the original model parameters are converted in $L_{eff}$ into $\varphi$-dependent classical effective parameters. In particular, the effective potential $U_{eff}(\varphi)$ in $L_{eff}$ has the form $U_{eff}=\frac{\lambda}{4\xi^{2}}M_{P}^{4}\cdot F(\varphi)\cdot\tanh^{4}\bigl(\frac{\sqrt{\xi}\varphi}{M_{P}}\bigr)$, where $F(\varphi)$ is a smooth function equal to $F(\varphi)\approx\frac{1}{2}$ for $\varphi>\sqrt{6}M_{P}$. It is fundamentally important that the constant $\xi$ of non-minimal coupling to the scalar curvature can be chosen small. If $\xi=1/6$, then to ensure agreement with CMB observational data, the Higgs field self-coupling parameter $\lambda$ in the original action must be of the order of $\sim 10^{-11}$. During cosmological evolution after the end of inflation, the decrease of $\varphi$ leads to a change in the sign of the effective Higgs mass term in $L_{eff}$. This TMSM effect provides an answer to the mystery of the Higgs potential structure and leads to spontaneous symmetry breaking. As $\varphi$ approaches VEV, the scalar function $\zeta(\varphi)$ changes in such a way that the classical TMT-effective self-coupling parameter $\lambda(\zeta(\varphi))$ increases by 10 orders of magnitude compared to $\lambda$, which is necessary for the implementation of the GWS theory. Applying the model to the very beginning of the classical evolution of the Universe shows that under certain initial conditions, cosmological dynamics can begin with a "pathological" and even phantom regime preceding inflation. However, if evolution begins with normal dynamics, then it proceeds only as inflation, and the problem of initial conditions for the onset of inflation does not arise. The fermion preheating model is described as a preliminary study of preheatig after inflation. Mathematical and physical arguments in favor of using the TMT are presented.
@@ 1 Introduction
To compare all existing Higgs inflation models with the model proposed in this paper, we begin with a brief description of the main steps of their construction at the tree level. The electroweak SM at the tree level is defined in Minkowski space. To use the SM in cosmology there is a need to reformulate it in a general covariant form to allow gravity to be included in the model. The next steps are to add a nonminimal coupling of the Higgs field to the scalar curvature, followed by the transition to the Einstein frame in the original action and a redefinition of the Higgs field leading to a Lagrangian with a canonically normalized field. This combination of particle physics and gravity, which we will call SM+Gravity, is the basis for how the Higgs sector of the electroweak SM is commonly used to study cosmology at energies well above the electroweak energy scale. However, the resulting models are suitable for describing inflation in accordance with the CMB data [20], [21] only if the constant of non-minimal coupling to scalar curvature is unnaturally huge: depending on the type of model, metric or Palatini, the coupling constant varies in the range $10^{4}\lesssim\xi\lesssim 10^{8}$. Nevertheless, apart from some effects due to nonminimal coupling (see, for example, review [10]), in the SM+Gravity approach to Higgs inflation models at the tree level, the question of finding a reasonable link between low-energy particle physics and physics at the inflation energy scale remains open. Moreover, the need to take into account quantum corrections runs into problems associated with the non-renormalizability of the resulting models, which further increases doubts about the validity of the SM+Gravity approach.
In the present paper, the Higgs inflation model is formulated in the framework of the Two-Measure Theory (TMT), which radically changes the approach to combining gravity with particle physics. Due to these changes, a formulation of the SM that takes into account the structure and procedure of TMT (and which we will refer to as Two-Measure Standard Model (TMSM)) can successfully overcome the above-mentioned problems of Higgs inflation.
@@
  • The main result of the present paper is the demonstration that the Higgs+gravity sector of TMSM with a small constant $\xi$ of non-minimal coupling to the scalar curvature allows one to construct a model of Higgs inflation that agrees well with observational CMB data. At the same time, despite the seemingly obvious expectations, there is no contradiction between this result and the fact that the TMSM also gives the correct values of the SM physical quantities at accelerator energies. Moreover, the TMSM provides a natural explanation for the mystery of the appearance of a mass term with the "wrong" sign in the Higgs potential as an effect arising during cosmological evolution due to a special property of the TMSM associated with the presence of two volume measures in the original action.
@@
are well known as the constraints neeeded for the onset of inflation. The initial understanding of this issue was developed in the first models of chaotic inflation [23], [24]. However, later, the need to implement flat potential models [25], [26], [27], [28], [29], [2], in which inflation predictions agree well with the SMB observational data, turned the question of initial conditions into a problem. The seriousness and urgency of this problem became clear from the discussion in papers [30]-[34] . A completely new approach to the problem of initial conditions for inflation based on the TMT was proposed in ref. [35], where it was shown that the solution to the problem is closely related to the need to answer the question of what happened before inflation. Within the framework of the TMT model of ref. [35], where the inflaton field was considered independent of the known matter fields, the problem of initial conditions for inflation turned out to be related to the Borde-Guth-Vilenkin (BGV) theorem [36]-[39]. In the present paper, in which the inflaton field is identified with the Higgs field, we find that not only are the initial conditions (1.1) for inflation guaranteed to be satisfied, but also that an earlier, rather exotic stage of evolution prior to inflation is also possible.
@@
The organization of the paper is as follows. In section 2, after defining the primordial action [^2] In ”conventional” alternative gravity theories, the original action differs from the Einstein-Hilbert action only in the form of the Lagrangian, for example, by the presence of a non-minimal coupling with curvature. The transformation to the Einstein frame, which in conventional theories is performed in the original action, simply changes the set of variables used to describe the theory, but does not change the theory itself. In TMT, the main difference is that the volume measure contains $\Upsilon$, although a modification of the Lagrangian is also possible. Therefore, the original TMT action differs from the Einstein-Hilbert action in the form of the Lagrangian density, and to bring the TMT action to the Einstein frame, the conformal transformation must exclude from the gravitational term of the action not only the non-minimal coupling, but also $\Upsilon$. In a theory obtained in this way, the constraint determining $\zeta$ cannot arise, that is, by acting in this way we would be dealing with a different theory. It is therefore incorrect to use the term ”Jordan frame” in the usual sense for the set of variables in the original action of TMT. To emphasize this distinction, we will use the terms ”primordial variables”, ”primordial model parameters”, ”primordial Lagrangian” and ”primordial action” instead of ”variables in the original frame”, ”original model parameters”, ”original Lagrangian” and ”original action”. and justifying the need to add a new type of vacuum-like term related to the volume measure density $\Upsilon$, we perform all the steps required by the TMT procedure. Having obtained the equations, we find that their consistency requires that the scalar $\zeta$ satisfy an algebraic constraint describing the dependence of $\zeta$ on the matter fields. By representing the equations in the Einstein frame and using the constraint, we arrive at the TMT-effective equations, potential, Lagrangian, and action [^3] Note that this use of the term ”effective” refers to the classical description and, for example, is in no way connected with the usual use of the term ”effective potential” calculated taking into account quantum corrections. for the Higgs sector of the TMSM (i.e. in the simplified TMSM, where all contributions of gauge fields and fermions are omitted). At this point, the reader wishing to understand the more fundamental reasons for the choice of the primordial action is advised to read Appendix A. Analysis of the obtained (classical) TMT-effective action shows that it is applicable in the entire admissible energy spectrum: from the energies of the electroweak SM to the scale of inflation. In section 3, after choosing the model parameters that provide the required flatness of the TMT effective potential for values of the canonically normalized scalar field $\varphi>6M_{P}$, we study a simplified model showing its agreement with the observational CMB data. The need to consider a simplified model (which will be justified in section 6) is due to the fact that the found TMT-effective action has a structure reminiscent of the K-essence models commonly used to describe the late time cosmic acceleration [40]-[43]. Important details of the shape of the TMT-effective potential and the behavior of the function $\zeta(\varphi)$ during the entire cosmological evolution are analyzed in section 4. This allows us to understand the role of cosmological evolution in the appearance of the negative mass term in the Higgs potential. Section 5, which is partly related to the paper [35], focuses on the effects dictated by the need to ensure that cosmological solutions (and their initial conditions) must keep the $\zeta$ function positive. In section 6 we study the influence of the K-essence type structure on the earliest stages of classical evolution of the Universe. We show that the K-essence type structure may be responsible for the existence of the earliest stages of evolution, which have "pathological" dynamics. However, if evolution starts with "normal" dynamics, then it proceeds only as inflation, and the problem of initial conditions for the onset of inflation does not arise. Finally, it is shown that the K-essence type structure does not change the results of the simplified version of the model obtained in section 3, as far as the slow-roll inflation stage is concerned. Section 7 presents perspectives for further developments, including a summary of the main results of TMSM in ref. [22] and key findings of the preliminary preheating study. In the Discussion section 8 we focus on fundamentally important features of TMSM that provide the main results of the paper. Appendix A is devoted to the rationale for the TMT approach developed in this paper. In the Appendix B it is shown that in the interval of $6M_{P}<\varphi<14M_{P}$, where the TMT-effective potential is flat and the slow-roll approximation is valid, the scalar $\zeta$ can vary within the range $0<\zeta\lesssim 5\cdot 10^{-3}$.
@@ 2.1 Action
Except for section 7, in this paper we will study the sector of TMSM with the following primordial action
$S_{primordial}=S_{gr}+S_{H}+S_{vac},$ (2.1)
where
$S_{gr}=\int d^{4}x(\sqrt{-g}+\Upsilon)\left(-\frac{M_{P}^{2}}{2}\right)\Biggl(1+\xi\frac{2|H|^{2}}{M_{P}^{2}}\Biggr)R(\Gamma,g)$ (2.2)
$S_{H}=\int d^{4}x\left[(b_{k}\sqrt{-g}-\Upsilon)g^{\alpha\beta}\left(\partial_{\alpha}H\right)^{{\dagger}}\partial_{\beta}H-(b_{p}\sqrt{-g}+\Upsilon)\lambda|H|^{4}-(b_{p}\sqrt{-g}-\Upsilon)m^{2}|H|^{2}\right].$ (2.3)
@@
Here $M_{P}$ is the reduced Planck mass and $H$ is the Higgs isodoublet. We will proceed in Palatini’s formulation, where $\Gamma$ denotes the affine connection, and $R(\Gamma,g)=g^{\mu\nu}R_{\mu\nu}(\Gamma)$, $R_{\mu\nu}(\Gamma)=R^{\lambda}_{\mu\nu\lambda}(\Gamma)$ and $R^{\lambda}_{\mu\nu\sigma}(\Gamma)\equiv\Gamma^{\lambda}_{\mu\nu,\sigma}+\Gamma^{\lambda}_{\gamma\sigma}\Gamma^{\gamma}_{\mu\nu}-(\nu\leftrightarrow\sigma)$. Following what was said in the Introduction, all contributions from gauge fields and fermions are omitted. In our notations, in a theory with only the volume element $\sqrt{-g}d^{4}x$, the model of massless scalar field non-minimally coupled to scalar curvature would be conformally invariant if the parameter $\xi$ were equal to $\xi=-\frac{1}{6}$. The model parameters $b_{k}>0$, $b_{p}>0$, $\lambda>0$ and $m^{2}>0$ are chosen to be positive. If the first term in the vacuum-like contributions (2.4) was present in Einstein’s GR, $V_{1}$ would be a cosmological constant. The term with $V_{2}$ was first introduced in Ref. [46]. The first reason for adding the term with $V_{2}$ is that the term $\propto\Upsilon=\zeta(x)\sqrt{-g}$, which can be expected as the contribution of quantum-gravitational effects to a vacuum-like action, does not contribute to the equations of motion. Therefore, the next in powers of $\zeta$ vacuum-like term can be of the form $\propto\zeta^{2}\sqrt{-g}=\Upsilon^{2}/\sqrt{-g}$. There are also more pragmatic reasons. It turns out that thanks to this term, the corresponding TMT-effective potential and the TMT-effective action acquire fundamentally new important properties without which the goals of this paper could not be achieved.
Given that the FLRW Universe is a model obtained by cosmological averaging, we can consider the corresponding averaging of the Higgs doublet
$\langle H(x)\rangle_{cosm.av.}=\frac{1}{\sqrt{2}}\left(\begin{matrix}0\\ \phi(t)\end{matrix}\right),$ (2.5)
and the classical homogeneous scalar field $\phi(t)$ is the only non-zero component of $\langle H(x)\rangle_{cosm.av.}$. In the model studied in this paper, $\phi(t)$ plays the role of the inflaton.
@@ 2.2 Equations of motion in the Einstein frame
$\frac{M_{P}^{2}}{2}\Omega R(\Gamma,g)-\frac{1}{2}g^{\alpha\beta}\phi_{,\alpha}\phi_{,\beta}-\frac{\lambda}{4}\phi^{4}+\frac{m^{2}}{2}\phi^{2}-2\zeta V_{2}={\mathcal{M}}$ (2.10)
@@
$(1+\zeta)\left(-\frac{M_{P}^{2}}{2}\right)\Omega R_{\mu\nu}(\Gamma)+(b_{k}-\zeta)\frac{1}{2}\phi_{,\mu}\phi_{,\nu}$
@@
Using the trace of eq.(2.11) one can eliminate the term $\frac{1}{2}M_{P}^{2}\Omega R(\Gamma,g)$ from eq.(2.10). As a result, we find that for these equations to be consistent, it is necessary that the scalar function $\zeta(x)$ satisfy the following relation
@@
[…]
@@
$\zeta=\zeta(\phi,X_{\phi};\mathcal{M})=\frac{\mathcal{M}-2V_{1}-(2b_{p}-1)\frac{\lambda}{4}\phi^{4}-(2b_{p}+1)\frac{m^{2}}{2}\phi^{2}+(1+b_{k})\Omega X_{\phi}}{\mathcal{M}-2V_{2}+\frac{\lambda}{4}\phi^{4}-\frac{m^{2}}{2}\phi^{2}-(1+b_{k})\Omega X_{\phi}}.$ (2.19)
@@
In the Higgs inflation models in the conventional theory (both in the metric and in the Palatini formulation), the transition to the Einstein frame performed directly in the original action completely eliminates the non-minimal coupling of the scalar field to the scalar curvature. In the TMSM under study, as in all TMT models, the transition to the Einstein frame (2.15) is carried out in the equations of motion and we notice that in the $\phi$-field equation (2.20) presented in the Einstein frame using (2.15), the term of the non-minimal coupling $\propto\xi R(\tilde{g})\phi$ is preserved. After passing to the Einstein frame in eq.(2.20), the $\phi$-field equation takes the following form
$\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}\Bigg(\frac{b_{k}-\zeta}{(1+\zeta)\Omega}\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}\phi\Bigg)+\frac{\xi}{\Omega^{2}}R(\tilde{g})\phi+\frac{b_{p}+\zeta}{(1+\zeta)^{2}\Omega^{2}}\lambda\phi^{3}+\frac{b_{p}-\zeta}{(1+\zeta)^{2}\Omega^{2}}m^{2}\phi=0,$ (2.21)
where R(g~)R($\tilde${g}) is determined from the Einstein equations (2.16) with the TMT-effective energy-momentum tensor (2.17), and $\zeta$ is given by the constraint (2.19). The procedure performed means that we are dealing with an explicitly formulated self-consistent system of the Higgs field and gravity.